### v1

@@ 3.2 A novel autoregressive Poisson lemma
The cornerstone of the design of the Stein’s Unbiased Risk Estimate is the seminal Stein’s lemma [60], turning an expectation explicitly involving ground truth, into another one in which the explicit dependency is completely removed. Under the Poisson noise model (11), the standard Stein’s lemma cannot be used. Further, due to the dependency of the memory term in past observations stated in Equation (7), none of the Stein’s lemmas generalized to Poisson noise [20, 42, 39] apply. The major challenge in designing oracles adapted to the driven autoregressive Poisson model is thus to derive a new autoregressive Poisson Stein’s lemma counterpart.
@@
is summable with respect to the driven autoregressive Poisson distribution obtained plugging the driven autoregression model (7) into the Poisson distribution (11)
@@
Let $\boldsymbol{\mathsf{Y}}=(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{T})$ observations following the driven autoregressive model (7) with ground truth time-varying reproduction coefficient $\overline{\boldsymbol{\mathsf{X}}}=(\overline{\mathsf{X}}_{1},\ldots,\overline{\mathsf{X}}_{T})\in\mathbb{R}_{+}^{T}$ and memory functions $\Psi_{t}$ satisfying Assumption 2, corrupted by scaled Poisson noise (11) of time-varying scale parameter $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{T})\in\mathbb{R}_{+}^{T}$. Then, for $\Theta:\mathbb{R}^{T}\rightarrow\mathbb{R}$ satisfying Assumption 1 $\forall t\in\{1,\ldots,T\}$,
@@
Proof of lemma 1 is detailed in Appendix A. ∎
@@ 3.3 Autoregressive Poisson Unbiased Risk Estimators
$\begin{split}&\mathsf{APURE}^{\mathcal{P}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda}\,\lvert\,\boldsymbol{\alpha})=\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\rVert^{2}_{2}\\ &-2\sum_{t=1}^{T}\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}+\sum_{t=1}^{T}\left(\mathsf{Y}_{t}^{2}-\alpha_{t}\mathsf{Y}_{t}\right)\end{split}$ (33)
@@
Further assuming that $\forall t\in\{1,\ldots,T\},$ $\Psi_{t}(\boldsymbol{\mathsf{Y}})\neq 0$, define the data-dependent estimation risk estimate
$\mathsf{APURE}^{\mathcal{E}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda}\,\lvert\,\boldsymbol{\alpha})=\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\rVert^{2}_{2}$ (35)
$-2\sum_{t=1}^{T}\frac{\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}+\sum_{t=1}^{T}\left(\frac{\mathsf{Y}_{t}^{2}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})^{2}}-\frac{\alpha_{t}\mathsf{Y}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})^{2}}\right).$
@@
Proof of Theorem 1 is developed in Appendix B. ∎
@@ 3.4 Finite difference Monte Carlo estimators
Let $\boldsymbol{\mathsf{Y}}=(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{T})$ be observations satisfying the requirements of Lemma 1 and $\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})$ be a parametric estimator of $\overline{\boldsymbol{\mathsf{X}}}$ whose components satisfy Assumption 1 as stated in Theorem 1 and satisfying Assumption 3. Let $\boldsymbol{\zeta}\sim\mathcal{N}(\boldsymbol{0},\textbf{I})$ a zero-mean Gaussian vector with covariance matrix the identity in dimension $T$. Define the data-dependent Finite Difference Monte Carlo prediction risk estimate
$\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda}\,\lvert\,\boldsymbol{\alpha})$
$=\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\rVert^{2}_{2}-2\sum_{t=1}^{T}\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}$ (37)
$+2\left\langle\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}))\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}],\mathrm{diag}(\boldsymbol{\mathsf{Y}})\boldsymbol{\zeta}\right\rangle+\sum_{t=1}^{T}\left(\mathsf{Y}_{t}^{2}-\alpha_{t}\mathsf{Y}_{t}\right)$
where $\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}]$ denotes the differential of $\boldsymbol{\mathsf{Y}}\mapsto\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})$ at the current point $(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})$ applied to the random vector $\boldsymbol{\zeta}$. Then, $\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}$ is an asymptotically unbiased estimate of the prediction risk in the small scale parameters limit, that is
@@
Further assuming that $\forall t\in\{1,\ldots,T\},$ $\Psi_{t}(\boldsymbol{\mathsf{Y}})\neq 0$, define the data-dependent Finite Difference Monte Carlo estimation risk estimate
$\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda}\,\lvert\,\boldsymbol{\alpha})=\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\rVert^{2}_{2}-2\sum_{t=1}^{T}\frac{\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}$
$+2\left\langle\mathrm{diag}(\boldsymbol{\alpha}\centerdot/\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}))\,\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}],\mathrm{diag}(\boldsymbol{\mathsf{Y}})\boldsymbol{\zeta}\right\rangle$ (39)
$+\sum_{t=1}^{T}\left(\frac{\mathsf{Y}_{t}^{2}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})^{2}}-\frac{\alpha_{t}\mathsf{Y}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})^{2}}\right).$
Then, $\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}$ is an asymptotically unbiased estimate of the estimation risk in the small scale parameters limit, that is
@@
Proof of Theorem 2 is developed in Appendix C. ∎
@@
$\begin{split}\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}=\frac{1}{N}\sum_{n=1}^{N}\mathsf{APURE}_{\boldsymbol{\zeta}^{(n)}}^{\mathcal{E}}\\ \overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}=\frac{1}{N}\sum_{n=1}^{N}\mathsf{APURE}_{\boldsymbol{\zeta}^{(n)}}^{\mathcal{P}}\end{split}$ (41)
@@
$\begin{split}&\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathcal{E}(\widehat{\boldsymbol{\mathsf{X}}},\overline{\boldsymbol{\mathsf{X}}})\\ &\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathcal{P}(\widehat{\boldsymbol{\mathsf{X}}},\overline{\boldsymbol{\mathsf{X}}}).\end{split}$ (42)
@@ 4 Application to piecewise linear estimation
The purpose of this section is twofold: first, to assess the ability of the robustified Finite Difference Monte Carlo risk estimates deriving from Theorem 2 and Proposition 1, to approximate faithfully the true estimation and prediction risks; second, to demonstrate numerically that the hyperparameters selected by minimizing these risk estimates yields accurate estimates of the reproduction coefficients from observations following the driven autoregressive model with data-dependent Poisson noise (11). To that aim, intensive simulations are run on synthetic data generated according to (7) and (11).
@@ 4.1 Synthetic data
To prepare for the application to epidemiological indicator estimation, developed in Section 5, the ground truth reproduction coefficient $\overline{\boldsymbol{\mathsf{X}}}$ is designed piecewise linear [1, 49], imitating temporal evolution of the reproduction number of COVID-19 observed from real-world data [18]. All synthetic data in this section are of length $T=70$ and share the same ground truth, represented by the deep blue curve in Figure 1, bottom plot, alternating expansion and recession phases, as represented by the blue areas indicating exponential growth period characterized by $\overline{\mathsf{X}}_{t}>1$. The initial state is set according to the observed real COVID-19 infection counts during the imitated period at $\mathsf{Y}_{0}=3395$. Mimicking the epidemiological model proposed in [15] particularized to COVID-19 pandemic, the memory functions are chosen linear, as described in Example 1, with a constant memory horizon of $\tau=25$ and a sequence $\{\psi_{s}\}_{s=1}^{\tau}$ chosen as the daily discretization of the serial interval distribution of COVID-19 modeled as a Gamma distribution of mean $6.6$ days and standard deviation $3.5$ days [8, 55]. The scale parameter of the data-dependent Poisson noise is constant through time, that is $\forall t,\alpha_{t}=\alpha>0$. Seven values of $\alpha$ logarithmically spaced between $\alpha=10^{2}$, corresponding to low noise level, to a very high noise level of $\alpha=10^{5}$, are explored. Figure 2 provides in top row examples of synthetic observations with the same underlying ground truth reproduction coefficient, displayed in blue in the second row plots, for three scale parameter values, corresponding, from first to third columns, to low $\alpha=10^{2}$, medium $\alpha=10^{3}$ and high $\alpha=10^{4}$ noise levels.
@@
Figure 2: Estimation of the reproduction coefficient from synthetic driven autoregressive Poisson observations using the variational estimator (43) coupled with the oracle-based hyperparameter selection strategy (4). First row: One realization of synthetic observations drawn following the setup described in Section 4.1. Second row: Underlying ground truth reproduction coefficient in blue, and estimates obtained from the variational estimator (43) combined with the oracle-based regularization parameter selection strategy, for the two ground truth-dependent oracles $\mathcal{P}^{\circ}$ and $\mathcal{E}^{\circ}$ in red and green respectively, and for the two data-driven oracles $\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ and $\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ in pink and orange respectively. Third and fourth rows: Robustfied data-driven prediction and estimation risk estimates computed on a logarithmic grid of regularization parameters $\lambda$, accompanied with their Gaussian confidence regions computed from the $N=10$ Monte Carlo vector realizations, and exact prediction and estimation errors; vertical dashed lines indicate optimal hyperparameters selected for each oracle.

### v3

@@ 3.2 A novel autoregressive Poisson lemma
The cornerstone of the design of the Stein’s Unbiased Risk Estimate is the seminal Stein’s lemma [71], turning an expectation explicitly involving ground truth, into another one in which the explicit dependency is completely removed. Under the Poisson noise model (11), the standard Stein’s lemma cannot be used. Further, due to the dependency of the memory term in past observations stated in Equation (7), none of the Stein’s lemmas generalized to Poisson noise [23, 51, 48] apply. The major challenge in designing oracles adapted to the generalized time-varying autoregressive Poisson model is thus to derive a new autoregressive Poisson Stein’s lemma counterpart.
@@
is summable with respect to the autoregressive Poisson distribution obtained plugging the generalized time-varying autoregression model (7) into the Poisson distribution (11)
@@
Let $\boldsymbol{\mathsf{Y}}=(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{T})$ observations following the generalized time-varying autoregressive model (7) with ground truth time-varying reproduction coefficient $\overline{\boldsymbol{\mathsf{X}}}=(\overline{\mathsf{X}}_{1},\ldots,\overline{\mathsf{X}}_{T})\in\mathbb{R}_{+}^{T}$ and memory functions $\Psi_{t}$ satisfying Assumption 2, corrupted by scaled Poisson noise (11) of time-varying scale parameter $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{T})\in\mathbb{R}_{+}^{T}$. Then, for $\Theta:\mathbb{R}^{T}\rightarrow\mathbb{R}$ satisfying Assumption 1 $\forall t\in\{1,\ldots,T\}$,
@@
Proof of lemma 1 is detailed in A. ∎
@@ 3.3 Autoregressive Poisson Unbiased Risk Estimators
$\begin{split}\mathsf{APURE}^{\mathcal{P}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda}\,\lvert\,\boldsymbol{\alpha})=\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\rVert^{2}_{2}-2\sum_{t=1}^{T}\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}+\sum_{t=1}^{T}\left(\mathsf{Y}_{t}^{2}-\alpha_{t}\mathsf{Y}_{t}\right)\end{split}$ (31)
@@
Further, assuming that $\forall t\in\{1,\ldots,T\},$ $\Psi_{t}(\boldsymbol{\mathsf{Y}})\neq 0$, define the data-dependent estimation risk estimate
$\mathsf{APURE}^{\mathcal{E}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda}\,\lvert\,\boldsymbol{\alpha})=\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\rVert^{2}_{2}-2\sum_{t=1}^{T}\frac{\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}+\sum_{t=1}^{T}\left(\frac{\mathsf{Y}_{t}^{2}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})^{2}}-\frac{\alpha_{t}\mathsf{Y}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})^{2}}\right).$ (33)
@@
Proof of Theorem 1 is developed in B. ∎
@@ 3.4 Finite Difference Monte Carlo estimators
Let $\boldsymbol{\mathsf{Y}}=(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{T})$ be observations satisfying the requirements of Lemma 1 and $\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})$ be a parametric estimator of $\overline{\boldsymbol{\mathsf{X}}}$ whose components satisfy Assumption 1 as stated in Theorem 1 and satisfying Assumption 3. Let $\boldsymbol{\zeta}\sim\mathcal{N}(\boldsymbol{0},\textbf{I})$ a zero-mean Gaussian vector with covariance matrix the identity in dimension $T$. Define the data-dependent Finite Difference Monte Carlo prediction risk estimate $\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda}\,\lvert\,\boldsymbol{\alpha})=$
$\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\rVert^{2}_{2}-2\sum_{t=1}^{T}\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}+2\left\langle\mathrm{diag}(\boldsymbol{\alpha}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}))\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}],\mathrm{diag}(\boldsymbol{\mathsf{Y}})\boldsymbol{\zeta}\right\rangle+\sum_{t=1}^{T}\left(\mathsf{Y}_{t}^{2}-\alpha_{t}\mathsf{Y}_{t}\right)$ (35)
where $\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}]$ denotes the differential of $\boldsymbol{\mathsf{Y}}\mapsto\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})$ at the current point $(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})$ applied to the random vector $\boldsymbol{\zeta}$. Then, $\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}$ is an asymptotically unbiased estimate of the prediction risk in the small scale parameters limit:
@@
Further assuming that $\forall t\in\{1,\ldots,T\},$ $\Psi_{t}(\boldsymbol{\mathsf{Y}})\neq 0$, define the data-dependent Finite Difference Monte Carlo estimation risk estimate $\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda}\,\lvert\,\boldsymbol{\alpha})=$
$\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\rVert^{2}_{2}-2\sum_{t=1}^{T}\frac{\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}+2\left\langle\mathrm{diag}(\boldsymbol{\alpha}\centerdot/\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}))\,\partial_{\boldsymbol{\mathsf{Y}}}\widehat{\boldsymbol{\mathsf{X}}}[\boldsymbol{\zeta}],\mathrm{diag}(\boldsymbol{\mathsf{Y}})\boldsymbol{\zeta}\right\rangle+\sum_{t=1}^{T}\left(\frac{\mathsf{Y}_{t}^{2}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})^{2}}-\frac{\alpha_{t}\mathsf{Y}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})^{2}}\right).$ (37)
Then, $\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}$ is an asymptotically unbiased estimate of the estimation risk in the small scale parameters limit:
@@
Proof of Theorem 2 is developed in C. ∎
@@
$\begin{split}\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}=\frac{1}{N}\sum_{n=1}^{N}\mathsf{APURE}_{\boldsymbol{\zeta}^{(n)}}^{\mathcal{E}}\quad\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}=\frac{1}{N}\sum_{n=1}^{N}\mathsf{APURE}_{\boldsymbol{\zeta}^{(n)}}^{\mathcal{P}}\end{split}$ (39)
@@
$\begin{split}\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathcal{E}(\widehat{\boldsymbol{\mathsf{X}}},\overline{\boldsymbol{\mathsf{X}}})\quad\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathcal{P}(\widehat{\boldsymbol{\mathsf{X}}},\overline{\boldsymbol{\mathsf{X}}}).\end{split}$ (40)
@@ 4 Application to piecewise linear estimation
The purpose of this section is twofold: first, to assess the ability of the robustified Finite Difference Monte Carlo risk estimates deriving from Theorem 2 and Proposition 1, to approximate faithfully the true estimation and prediction risks; second, to demonstrate numerically that the hyperparameters selected by minimizing these risk estimates yields accurate estimates of the reproduction coefficients from observations following the generalized time-varying autoregressive model with data-dependent Poisson noise (11). To that aim, intensive simulations are run on synthetic data generated according to (7) and (11).
@@ 4.1 Synthetic data
To prepare for the application to epidemiological indicator estimation, developed in Section 5, the ground truth reproduction coefficient $\overline{\boldsymbol{\mathsf{X}}}$ is designed piecewise linear [1, 59], imitating temporal evolution of the reproduction number of COVID-19 observed from real-world data [21]. All synthetic data in this section are of length $T=70$ and share the same ground truth, represented by the deep blue curve in Figure 1, bottom plot, alternating expansion and recession phases, as represented by the blue areas indicating exponential growth period characterized by $\overline{\mathsf{X}}_{t}>1$. The initial state is set according to the observed real COVID-19 infection counts during the imitated period at $\mathsf{Y}_{0}=3395$. Mimicking the epidemiological model proposed in [17] particularized to COVID-19 pandemic, the memory functions are chosen linear, as described in Example 1, with a constant memory horizon of $\tau=25$ and a sequence $\{\psi_{s}\}_{s=1}^{\tau}$ chosen as the daily discretization of the serial interval distribution of COVID-19 modeled as a Gamma distribution of mean $6.6$ days and standard deviation $3.5$ days [9, 66]. The scale parameter of the data-dependent Poisson noise is constant through time, that is $\forall t,\alpha_{t}=\alpha>0$. Seven values of $\alpha$ logarithmically spaced between $\alpha=10^{2}$, corresponding to low noise level, to a very high noise level of $\alpha=10^{4.5}$, are explored. Figure 2 provides in top row examples of synthetic observations with the same underlying ground truth reproduction coefficient, displayed in blue in the second row plots, for three scale parameter values, corresponding, from first to third columns, to low $\alpha=10^{2}$, medium $\alpha=10^{3}$ and high $\alpha=10^{4}$ noise levels.
@@
Figure 2: Estimation of the reproduction coefficient from synthetic generalized time-varying autoregressive Poisson observations using the variational estimator (41) coupled with the oracle-based hyperparameter selection strategy (4). First row: One realization of synthetic observations drawn following the setup described in Section 4.1. Second row: Underlying ground truth reproduction coefficient in blue, and estimates obtained from the variational estimator (41) combined with the oracle-based regularization parameter selection strategy, for the two ground truth-dependent oracles $\mathcal{P}^{\circ}$ and $\mathcal{E}^{\circ}$ in red and green respectively, and for the two data-driven oracles $\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ and $\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ in pink and orange respectively. Third and fourth rows: Robustified data-driven prediction and estimation risk estimates computed on a logarithmic grid of regularization parameters $\lambda$, accompanied with their Gaussian confidence regions computed from the $N=10$ Monte Carlo vector realizations, and exact prediction and estimation errors; vertical dashed lines indicate optimal hyperparameters selected for each oracle.